The equation of the straight line passing through (1,1) and makes an angle 45 with the line 2x-y+7=0

Grade:11

1 Answers

Aditya Gupta
5 years ago
the equation will be
y – 1= m(x – 1), m being the slope
angle θ between two lines is given as
tanθ= mod((m1-m2)/(1+m1m2))
now, 2x-y+7=0 has slope 2
so, tan45=1= mod((m-2)/(1+2m))
or m-2= ±(1+2m)
or m= – 3, 1/3
so equation is 
y+3x=4 or 3y=x+2